OpenStudy (anonymous):

The length of a rectangle is 12x and the width is 4+ x. If the perimeter is 60 inches , then the value of x is: A=2 b=8 C=26

OpenStudy (imstuck):

What is the formula for a perimeter of a rectangle? Do you know?

OpenStudy (anonymous):

A=L * w rectangle

OpenStudy (imstuck):

that's correct for the area, but you are looking for the perimeter.

OpenStudy (anonymous):

i think that is the formulsa

OpenStudy (imstuck):

The formula for the perimeter of a rectangle is P = 2L + 2W. Does that look familiar?

OpenStudy (anonymous):

no

OpenStudy (imstuck):

Ok, well that's the formula you need. the perimeter is the measure along the OUTSIDE of the rectangle, while the area is the measure INSIDE the rectangle. Using the formula and the info we are given, sub in and solve for x. they tell you that the perimeter is 60, so P = 60. They also tell you that the length is 12x and the width is 4 + x. We need to fill those values into the formula, all of them, to solve for x. So P = 2L + 2W becomes 60 = 2(12x) + 2(4 + x). Do you know what to do with that, what steps come next?

OpenStudy (anonymous):

you combine like terms right?

OpenStudy (imstuck):

You have to do some multiplying first. You need to distribute the 2 into the 12x and into the (4 + x). Do you know how to do that?

OpenStudy (anonymous):

24x + 8+2x

OpenStudy (imstuck):

That's right! Now set that equal to 60, because 60 is your perimeter length and those are what you add together to get the perimeter. Set it equal to 60 and solve it for x. what do you get when you do that?

OpenStudy (anonymous):

24x+8+2x=60

OpenStudy (imstuck):

Ok, now combine like terms. Then get the x term on one side of the equals sign and the constant on the other side of the equals sign. How would you do that?

OpenStudy (anonymous):

I got 26x + 8=60?

OpenStudy (anonymous):

the answer is 2

OpenStudy (anonymous):

thanks @IMStuck

OpenStudy (imstuck):

Yes it is! Great! So obviously you had no problems with this? Very cool for you!

OpenStudy (anonymous):

Thank you so much :D :D :D :D :D :D :D :D